AÖF Soru Bankası
İST209U

Statıstıcs I (ENG)

Ünite Sınavı Soruları

  • 20 soru
  • Soru metinleri açık; şıklar ve cevap anahtarı uygulamada
  1. 1

    According to classical probability approach, what is the probability of obtaining a number 6 when you throw a six-sided fair die?

  2. 2

    I. For any event A⊆S, P(A)≥0. 

    II. For any event A, P(Ā) = 1 – P (A)

    III. For any two events A and B, P(A,B) = P(A) + P(B)

    Which of the probability axioms can be said to be true?

  3. 3

    A s­ix-si­ded fai­r dic­e has been thrown 3000 times and the occurrence of number 4 ­is 450. What i­s the empirical probability of obtaining a number 4 when you throw a s­ix-si­ded fai­r dic­e?

  4. 4

    For two sets A and B the probability of A is 0.15 and the probability of B is 0.25 if P(A|B) is equal to 0.30 what is the value of P(B|A)?

  5. 5

    According to classical probability, when a fair dice is rolled, what is the probability of obtaining a number more than 4?

  6. 6

    For two sets A and B the probability of B is 0.45 and P(A|B) is equal to 0.20 what is the value of P(A∩B)?

  7. 7

    There are 36 cards in a box with three different colors-6 yellow, 12 blue, 18 red. What is the probability of obtaining two yellow cards when randomly drawing two cards? 

  8. 8

    A company works with two supplier firms for the same raw material. The first firm's late delivery probability on a given day is 0.05, whereas for the second firm this probability is 0.07. In addition, both firms' late delivery probability on given day is 0.04. What is the late delivery probability of at least one of the firms to company on given day?

  9. 9

    In a study about employess, employees are classified according to the department they work in and their preference of training coursec. If there are 12 different departments and 7 different courses, in how many different ways can an employee be classified?

  10. 10

      What does this equation tell us?

  11. 11

    At a local district, there are three d­ifferent pizza restaurants denoted here by A, B, and C. It is known that 40% of all customers give an order from company A, whereas 35% give an order from company B, and 25% give an order from company C. It is also known that 10% of the motorcycles from company A, 20% of the motorcycles from company B, and 5% of the motorcycles from company C need to a checkup before the next pizza delivery. If a motorcycle returned to the company needs a checkup before the next pizza delivery, what is the probability that this motorcycle was owned by company B?

  12. 12

    We know that event B has occurred and we are interested in finding the probability of event A. In other words, we are interested in finding the probability of A knowing that event B has occurred. Which of the following formula denotes the probability? We know that event B has occurred and we are interested in finding the probability of event A. In other words, we are interested in finding the probability of A knowing that event B has occurred. Which of the following formula denotes the probability?

  13. 13

    I. The complement of an event A is the set of all basic outcomes in S that do not belong to A.

    II. The union of events A and B is the set of all elementary outcomes that belong to both sets. 

    III. The intersection of events A and B is the set of all elementary outcomes that belong to at least one of the sets A and B.

    For A and B, which are any two events in a random experiment with sample space S, which of the statements are true?

  14. 14

    If there were two dices thrown at the same time what would their total number of possible outcomes be?

  15. 15

    According to empirical probability, which of the following statements is not true?

  16. 16

    2 models will be randomly selected to walk in a fashion show in France. There are 2 models from an Italian agency and 3 of them work for a Turkish agency.  If the event “one Turkish model Y and one Italian model X is chosen” is denoted by E, which one below is the listing the elements of E?

     
  17. 17

    Consider the random experiment of tossing a fair coin until a tail (T) and a head (H) show up once. Describe the sample space of this random experiment ?

  18. 18

    In how many different ways can the letters in P R O B A B I L I T Y be arranged?

  19. 19

    Which of the followings cannot be given as an example for random experiments?

  20. 20

    According to classical probability, when rolling a fair dice, what is the probability of obtaining the number 5?

Şıkları ve doğru cevapları görStatıstıcs I (ENG) dersinin tüm sınavları, cevap anahtarları, süreli mod ve yanlış analizi uygulamada.Uygulamada aç